Find the values of p and q if (x-1) and (x+2) are both factors px*3 +4x*2+qx -6
1
Expert's answer
2019-08-22T11:53:05-0400

Using the factor theorem, if "f(x)=px^3+4x^2+qx-6" has factors "(x-1)" and "(x+2)", then

"f(1)=0, f(-2)=0".

"\begin{cases} p\cdot 1^3 +4\cdot 1^2 +q\cdot 1-6=0 \\ p\cdot (-2)^3 +4\cdot (-2)^2 +q\cdot (-2)-6=0 \end{cases}"

"\begin{cases} p+4 +q-6=0 \\ -8p +16 -2q-6=0 \end{cases}"

"\begin{cases} q=2-p \\ -8p -2q=-10 \end{cases}"

"\begin{cases} q=2-p \\ 4p+q=5 \end{cases}"

"\begin{cases} q=2-p \\ 4p+2-p=5 \end{cases}"

"\begin{cases} q=2-p \\ 3p=3 \end{cases}"

"\begin{cases} p=1 \\ q=1 \end{cases}"

Answer: "p=1,q=1"